how to solve it ?

.47. The form of a sound wave travelling through air is S(x,t) = S cos (kx + 3000t +の, where x is in meters and t in seconds. What is the shortest time interval that any air molecule takes along the path to move between displacements SS/ and SSm/3? (Ans: 0.23 ms)
Answer #1

comparing with standard equation

T = 2 pi /3000 seconds

Sm/3 = Sm cos@

cos@ = 1/3

@ = 70.53 deg

similarly

-Sm/3 = Smcos@

@ = 109.47 deg

d@ = 109.47 - 70.53 = 38.94 deg = 0.216pi radians

As in time T, wave covers 2 pi

2 pi / 0.216 pi = 2 pi / 3000 / dt

4.622 = 2.093 x10^-3/ dt

dt = 0.45 s